The length of a rectangle is three times it’s width if the perimeter is 72 calculate the width and the area of the rectangle
Answer by math_helper(2461)


L = 3W
P = 2L + 2W = 2(3W) + 2W = 8W
P = 72
8W = 72 --> W=9 units. L=3(9)=27 units, and area = 9*27 = 243 sq units.